Jim began a 192​-mile bicycle trip to build up stamina for a triathlete competition.​ Unfortunately, his bicycle chain​ broke, so he finished the trip walking. The whole trip took 6 hours. If Jim walks at a rate of 5 miles per hour and rides at 41 miles per​ hour, find the amount of time he spent on the bicycle.
Found 2 solutions by ankor@dixie-net.com, josgarithmetic:
Answer by ankor@dixie-net.com(22740)

Jim began a 192​-mile bicycle trip to build up stamina for a triathlete competition.​
Unfortunately, his bicycle chain​ broke, so he finished the trip walking.
The whole trip took 6 hours.
If Jim walks at a rate of 5 miles per hour and rides at 41 miles per​ hour, find the amount of time he spent on the bicycle.
:
let b = time on the bike
whole trip was 6 hrs, therefore
(6-b) = time walking
:
write a dist equation, dist = speed * time
bike dist + walk dist = 192 mi
41b + 5(6-b) = 192
41b + 30 - 5b = 192
41b - 5b = 192 - 30
36b = 162
b = 4.5 hrs on the bike
:
:
Check this, find time walking: 6 - 4.5 = 1.5 hrs
find the sum of the distance
41 * 4.5 = 184.5 mi
5 * 1.5 = 7.5 mi
--------------------
total dist: 192

Answer by josgarithmetic(39616)

               SPEEDS        TIME       DISTANCES

RIDING         41             x           41x

WALKING         5            6-x          5(6-x)

Totals                        6            192  

Sum of the distances give an equation.....